Sol Summary – Nov 6th

Summary Title: Water from Stone
Author’s name: Guy Murphy

Mission Status: Active.

Sol Activity Summary: Today Andrew succeeded in extracting water from gypsum crystals in the Science Dome for the first time this mission. Collection of samples for the micro-meteorite study continued during the morning EVA. Various Hab maintenance tasks were undertaken. 3 cooked meals were prepared, with Sandy baking a second loaf of bread.
Look Ahead Plan: We hope to complete the micro-meteorite sample collection from the active sample site by Friday. We are also planning the checkout process for our crew members who are leaving.

Anomalies in work: Broke sim briefly at around 4 pm to speak with 2 unexpected visitors in the hab environs, who were in fact lost.

Weather: Clear sunny day, no precipitation.

Crew Physical Status: Crew continues to be in good health.

EVA: Successfully completed.

Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Greenhab Report, Operations Reports, Photos.

Support Requested: None.

Science Report – Nov 6th

Crew 214 Science Report 6 November 2019
SOL: 10
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The seventh scientific EVA for the mission saw us return to Tank Wash #2 to collect 23 samples from the micrometeorite grid. Four of these had been compromised and it remains to be seen how this will affect the study. 64% of the plots have now been sampled.

With the disappointing failure of the regolith heater prior to arrival at MDRS, a modified plan to demonstrate water extraction was devised. Instead of an in situ test of the device, the gypsum and gypsiferous soil samples collected yesterday would undergo the process in the Science Lab. After documenting the samples, 6 small subsamples of ground gypsum were heated. Within minutes of each subsample being heated, water vapour was observed and video recorded. Further testing of the gypsiferous soil will be attempted tomorrow.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from yesterday’s EVA were processed and experimental procedures commenced.

Closed Loop Food Waste Study: The day’s waste for Sol 10 is: 51 grams; consisting of 175 calories, 4g fat, 2g saturated fat, 328mg sodium, 26g carbohydrates, 3g fibre, 8g sugar, 10g protein.

EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.

EVA Suit Maintenance: Nominal

Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

EVA Report – Nov 6th

Crew 214 EVA Report 06-11-2019
EVA #9
Author: Guy Murphy

Purpose of EVA: Geological Sampling

Start time: 0930
End time: 1200

Narrative: The ninth EVA for science saw us returning to the Tank Wash location #2 at 518477E 4253767N where we continued sampling the 1m x 1m plots for micrometeorites. The EVA allowed us to collect 23 samples. We anticipate 2 x normal length or possibly 1 x longer duration EVAs will complete this sampling program. We have now completed 64% of the grid, which included 4 grid squared compromised by footprints.

Destination: Drove from the hab along Cow Dung Road to Tank Wash and then returned to the hab.
Coordinates (use UTM NAD27 CONUS): 518477E 4253767N
Participants: Guy Murphy, Sandy Dance
Road(s) and routes per MDRS Map: Cow Dung Road to Tank Wash
Mode of travel: Spirit Rover, Curiosity Rover

Operations Report – Nov 6th

Crew 214 Operations Report 6-11-2019
SOL: 10
Name of person filing report: Andrew Wheeler

Non-nominal systems: Opportunity, Suit 10

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run
Hours run: 12h 40m
From what time last night: 1920
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 68%
Diesel Reading – 60%
Station Propane Reading – 80%
Ethanol Free Gasoline: N/A

Water (loft tank): 37 gallons
Water Meter: 1455889 units
Water (static tank): 223 gallons
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons

Toilet tank emptied: no

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 108.4
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 66%
Currently charging: yes

Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Curiosity rover used: yes
Hours: 114.3
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 74%
Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? N/A
CrewCar used and why, where? N/A

General notes and comments: The generator failed to recharge the batteries during the night.

Summary of internet: Functional
Summary of suits and radios: Nominal
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. 6 bags of soil obtained, which will be delivered to the Greenhab tomorrow.
Summary of Science Dome operations: All systems nominal. Sample processing following EVA was carried out and heating demonstration of hydrated minerals carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: N/A

Greenhab Report – Nov 6th

Crew 214 GreenHab Report 6-Nov-2019
GreenHab Officer: Guy Murphy

Environmental control: Door left open for ventilation during the afternoon then reclosed.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: Existing plantings watered with 24 litres (3x blue watering can).
Daily water usage for research and/or other purposes: N/A
Water in Blue Tank _____ gallons
Time(s) of watering for crops: 3 pm.
Changes to crops: N/A

Narrative: Apart from watering the plants, the only work undertaken in the Greenhab today was trimming of dead material from the plants including dead leaves, stalks.

Established plants introduced to Greenhab 04/11/2019

[3x] strawberry plants (Everbearing), in a medium metal planter
[2x] small Aloe Vera, in a medium metal planter, frost damaged
[1x] spicy orange thyme (thymus fragantissimus), small terra-cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra-cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris
Harvest: None
Support/supplies needed: N/A

Journalist Report – Nov 6th

Wed 6 Nov Sol 10
by Sandy Dance

Another day, another EVA! But we are not so blasé. T

his morning Guy and I went out to the micrometeorite grid to gather more magnetic particles. We motor along the road in our space suits and buggies, turn a corner and there is a large collection of vehicles. Oh no, our patch has been invaded!

Closer inspection reveals a film crew with all you would expect: catering, equipment, props. We wave as we pass, and get multiply photographed in return. They seemed to take the sight of Martians motoring by in their stride.

The micrometeorite gathering was a success, having done a couple more rows of the grid, dealing with the footprints left by an unknown interloper the other day by noting which squares were contaminated and carrying on.

Then back at the Science Dome Andrew begins analysing yesterday’s gypsum by heating it and seeing the water condensation on a beaker, then finding the weight difference: 0.4g of water from 17g of rock.


Later that afternoon we spot a couple of strangers walking on the hill near the Outpost. After failing to raise Shannon on the radio, we break sim and investigate. Turns out they were heading to the film crew up the road but got confused by our rather SciFi setup here. The film, they tell us, features futuristic Native Americans in a post-apocalyptic dystopia.

Last night we were successful spotting Saturn with the Celestron telescope: great sightseeing the rings!

Copyright © The Mars Society. All rights reserved. | Main Site